Hi i have a problem with finding the points of intersection of two polar equations. Can you please help me?
Find the points of intersection of two 4-petal roses given by: r=sin2theta and r=cos2theta.

Answer by Alan3354(69443)

Find the points of intersection of two 4-petal roses given by: r=sin2theta and r=cos2theta.
------------------
Since they both = r,
sin(2t) = cos(2t) = sqrt(1 - sin^2(2t))
sin^2(2t) = 1 - sin^2(2t)
sin^2(2t) = 1/2
sin(2t) = sqrt(2)/2 (plus and minus)
--> 2t = 45 + n*90 deg
t = 22.5 + n*45 degs